A 4.112-g tablet, containing quinine was dissolved in 0.10 M HCl to give 500 mL. Dilution of a 10.0-mL aliquot to 100 mL yielded a reading for fluorescence intensity (at 347.5 nm) of 320 on an arbitrary scale. A second 10.0-mL aliquot was mixed with 10.0 mL of 100-ppm quinine solution before dilution to 100 mL. The fluorescence intensity of this solution was 433. Calculate the percentage of quinine in the tablet.
Answer #1

Answers Quinime is a 1.664 g Antimalarial tablet was. dissolved in suficient o lom Hel to give 500ml of solution A 15ml Aliqulome (oolol) Aliquot @ 8.224 9 tablet/L = 0.08224 9 tablet in Aliquot. 0.08224 9 tablet into looml solution = 0.82249 tabletNow you have o.074 g quinimell, and that Come from a solution of 0.82249 tablet /L. So You have 0.074 9 quimime 10.82249 tabl
Answer #2

Let the quinine present in tablet = m

then m is dissolved in 500 ml and 10 ml of this taken which means ( m/500) x 10 = 0.02 m present in that 10 ml.

Now this gives reading 320.

Now 2nd aliquot of 10 ml ( containing 0.02m quinine) taken and 100 ppm quinine added.

100 ppm means 100 g per 10^ 6 ml solution. Since 10 ml of 100 ppm used, quinine mass =( 10 ml x 100 g / 10^ 6ml) = 0.001 g

now total quinine present when both added = 0.02m + 0.001 g

This gives reading 433

Absorbance = e x C x l where C is concentrtaion

A2/A1 = C2 /C1

433 / 320 = ( 0.02m+0.001) / ( 0.02m)

1.353 ( 0.02m) =0.02m + 0.001

0.027m =0.02m + 0.001

m = 0.142g

% of quinine = ( mass of quinine / mass of tablet) x 100

= ( 0.142g / 4.122g) x 100

= 3.44 %